OpenStudy (anonymous):

On the first day of school, the percentage of boys in a particular class is 60%. During the school year six girls move away and are replaced in the class by six boys; this makes the class roster 75% boys. Find the number of boys and girls in the class on the first day of school

OpenStudy (anonymous):

24 boys 16 girls

OpenStudy (anonymous):

6 boys = 15% 40 students = 100%

OpenStudy (anonymous):

Thanks sorry i stepped away! How did you set the equation up ??

OpenStudy (id21):

x... girls at 1st day. y... boys at 1st day. \[\frac{y}{x+y}=0.6\] ... boys percentage at 1st day. \[\frac{y+6}{x-6+y+6}=\ 0.75\] ...boys percentage after replacing. So you must solve this equation system. \[\frac{y}{x+y}=\frac{3}{5}\] \[\frac{y+6}{x+y}=\frac{3}{4}\] Find x from 1st equation: \[x=\frac{5y-3y}{3}=\frac{2y}{3}\] Put it in 2nd equation to find y: \[\frac{y+6}{\frac{2y}{3}+y}=\frac{3}{4}\] \[y=24\] And find the x: \[x=\frac{2\times 24}{3}=16\] So answer is that at 1st day there was 16 girls and 24 boys in the class.